# 7B Fully Open Source Moxin-LLM – From Pretraining to GRPO-based Reinforcement Learning Enhancement

### v3

@@ 1 Introduction
  • For the pre-training, we release the Moxin Base 7B model, together with all training code, data, and checkpoints, aiming to make continuous commitments to fully open-source LLMs. Our base model performs competitively compared with SOTA pre-trained base models, with a moderate training cost about \$160,000.
  • To further improve the model performance, we post-train our base model with SFT and DPO methods, based on the open-source Tülu 3 framework. Two versions of instruction models are developed with two different open-source datasets (Tülu 3 and Infinity Instruct), respectively. Both models are available and the one finetuned with the Tülu 3 dataset is released as Moxin Instruct 7B, with superior performance in different evaluations.
  • To improve the reasoning capabilities of the model with CoT, we further finetune our instruct model with SFT based on reasoning data from Open-Thoughts and OpenR1-Math-220k, and then reinforcement learning (GRPO following DeepSeek R1). We develop two versions of reasoning models with two different open-source RL frameworks (DeepScaleR and AReaL) and their corresponding data, respectively. Both models are available and the one finetuned with DeepScaleR is released as Moxin Reasoning 7B, with superior performance in CoT evaluations. We demonstrate that RL such as GRPO can work effectively for small LLMs.
@@ 3.4 Training Configuration
We use Colossal-AI [98] as our training framework with acceleration techniques including optimized parallelism, memory optimization, and heterogeneous training methods. : 1) Optimized Parallelism: We combine model, data, and pipeline parallelism to scale training across GPUs. This hybrid parallelism strategy minimizes communication overhead and ensures efficient resource utilization. 2) Memory Optimization: Techniques such as ZeRO redundancy elimination, gradient checkpointing, and mixed-precision training (FP16) are employed to maximize the effective batch size and reduce memory consumption. 3) Heterogeneous Training Methods: the dynamic scheduling and asynchronous communication protocols are leveraged to further boost training throughput. These methods are provided through user-friendly APIs, requiring minimal code modifications. The training cost is around \$160,000.
@@ 6 Evaluation
We conducted comprehensive performance comparisons against leading language models of comparable scale, including Mistral-7B [6], LLaMA 2-7B [71], Gemma-7B [56], and Qwen v2-7B [11]. These models were selected based on their demonstrated excellence within the 7B or 8B category and represent diverse development approaches from various research organizations worldwide. To ensure a robust evaluation, we re-run all benchmarks with the same evaluation pipeline for fair comparisons. Specifically, we use lm-evaluation-harness [111] and opencompass [112] for evaluation.
@@ 6.2 Pre-Training Evaluation
We name the initial model as Moxin-7B-original, which presents the foundation model before fine-tuning on the training data of the evaluation datasets. After subsequent partial fine-tuning of Moxin-7B-original on the training data of the evaluation datasets, we developed Moxin-7B-finetuned, enabling direct assessment of how targeted fine-tuning affects model performance. We release our Moxin-7B-finetuned model as Moxin 7B Base model.
@@ 6.2.1 Zero-Shot Evaluation
We report the result of base models for zero-shot evaluation in Table 6.2.1. The tasks are listed below. After training with the training data of evaluation tasks, our Moxin-7B-finetuned can achieve superior performance compared with state-of-the-art (SOTA) baselines. This significant increase from the base model demonstrates the effectiveness of our fine-tuning approach. The improved performance is particularly notable on complex reasoning tasks like PIQA, where the score increased from 78.07% to 82.24%, matching or exceeding several leading models. Consequently, our models emerge as an excellent candidate for real-world applications.
@@
[t] Performance comparison for various models in zero-shot evaluation. Models HellaSwag WinoGrade PIQA ARC-E ARC-C Ave Mistral - 7B 80.39 73.4 82.15 78.28 52.22 73.29 LLaMA 2 - 7B 75.99 69.06 79.11 74.54 46.42 69.02 LLaMA 2 - 13B 79.37 72.22 80.52 77.4 49.06 71.71 LLaMA 3.1 - 8B 78.92 74.19 81.12 81.06 53.67 73.79 gemma - 7b 80.45 73.72 80.9 79.97 54.1 73.83 Qwen v2 - 7B 78.9 72.38 79.98 74.71 50.09 71.21 internlm2.5 - 7b 79.14 77.9 80.52 76.16 51.37 73.02 Baichuan2 - 7B 72.25 67.17 77.26 72.98 42.15 66.36 Yi-1.5-9B 77.86 73.01 80.74 79.04 55.03 73.14 deepseek - 7B 76.13 69.77 79.76 71.04 44.8 68.3 Moxin - 7B - original 72.06 66.31 78.07 71.47 48.15 67.21 Moxin - 7B - finetune 80.03 75.17 82.24 81.12 58.64 75.44
@@ 6.2.2 Few-Shot Evaluation
Table 6.2.2 presents our zero-shot evaluation results across multiple benchmark tasks. The tasks and their few-show settings are listed below. Thanks to its rigorous and high-quality training corpus, our model demonstrates a remarkable competitive edge in tasks that involve language understanding and knowledge application. Our Moxin-7B-original achieves superior performance than LLaMA2-7B in this scenario. After training with the training data of evaluation tasks, our Moxin-7B-finetuned can achieve competitive performance compared with SOTA baselines.
@@
[t] Performance comparison for various models in few-shot evaluation. Model ARC-C Hellaswag MMLU WinoGrade Ave Mistral - 7B 57.59 83.25 62.42 78.77 70.51 LLaMA 3.1 - 8B 54.61 81.95 65.16 77.35 69.77 LLaMA 3 - 8B 55.46 82.09 65.29 77.82 70.17 LLaMA 2 - 7B 49.74 78.94 45.89 74.27 62.21 Qwen 2 - 7B 57.68 80.76 70.42 77.43 71.57 gemma - 7B 56.48 82.31 63.02 78.3 70.03 internlm2.5 - 7B 54.78 79.7 68.17 80.9 70.89 Baichuan2 - 7B 47.87 73.89 54.13 70.8 61.67 Yi-1.5-9B 58.36 80.36 69.54 77.53 71.48 Moxin - 7B - original 53.75 75.46 59.43 70.32 64.74 Moxin - 7B - finetuned 59.47 83.08 60.97 78.69 70.55
@@ 6.3 Post-Training Evaluation
In our post-training, with Tülu 3, we further fine-tune our base model with SFT and then DPO, leading to our SFT model and DPO model. We also adopt the Infinity Instruct dataset [63] to train another version of our instruct model based on our base model, and release this instruct model as Moxin-7B-II. We demonstrate the performance in Table 6.3 for zero-shot evaluation and Table 6.3 for few show evaluation. We also use the OLMES framework from Tulu 3 [62] to evaluate our model across a range of tasks and show the results in Table 6.3. We can observe that our DPO model can achieve comparable performance with other SOTA instruct models. We release our DPO model as Moxin 7B Instruct model.
@@ 6.4 CoT Evaluation
Based on our instruct model, we further finetune the model with reasoning data based on SFT and then train using reinforcement learning with GRPO following DeepSeek R1. Our final model is trained with DeepScaleR [67] and released as Moxin 7B Reasoning model. We also trained another version of the reasoning model with AReal [68] starting from our instruct model, with its name as Moxin - 7B - Reasoning - AReal. We evaluate our reasoning models on competition-level mathematics benchmarks, including AMC 2023, MATH-500, Minerva Math, and OlympiadBench. We report the Pass@1 accuracy. We use Qwen Math’s codebase for evaluation following [109, 110] and show the comparison in Table 6.4. We can observe that Moxin Reasoning model can achieve outstanding performance compared with baselines, demonstrating the effectiveness of RL for small LLMs such as 7B models. Our Moxin Reasoning model trained with DeepScaleR performs better than that trained with AReal.
[t] Performance comparison for various models on reasoning evaluation. Models/Datasets MATH 500 AMC Minerva Math OlympiadBench Ave Qwen2.5-Math-7B-Base 52.4 52.5 12.9 16.4 33.55 Qwen2.5-Math-7B-Base + 8K MATH SFT 54.6 22.5 32.7 19.6 32.35 Llama-3.1-70B-Instruct 64.6 30.1 35.3 31.9 40.48 Moxin - 7B - Reasoning - AReal 68.6 50 16.9 31.7 41.8 Moxin - 7B - Reasoning 68 57.5 16.9 30.4 43.2

### v4

@@ 1 Introduction
  • For the pre-training, we release the Moxin-7B-Base model, together with all training code, data, and checkpoints, aiming to make continuous commitments to fully open-source LLMs. Our base model performs competitively compared with SOTA pre-trained base models, with a moderate training cost about \$160,000.
  • To further improve the model performance, we post-train our base model with SFT and DPO methods, based on the open-source Tülu 3 framework. Two versions of instruction models are developed with two different open-source datasets (Tülu 3 and Infinity Instruct), respectively. Both models are available and the one finetuned with the Tülu 3 dataset is released as Moxin-7B-Instruct, with superior performance in different evaluations.
  • To improve the reasoning capabilities of the model with CoT, we further finetune our instruct model with SFT based on reasoning data from Open-Thoughts and OpenR1-Math-220k, and then reinforcement learning (GRPO following DeepSeek R1). We develop two versions of reasoning models with two different open-source RL frameworks (DeepScaleR and AReaL) and their corresponding data, respectively. Both models are available and the one finetuned with DeepScaleR is released as Moxin-7B-Reasoning, with superior performance in CoT evaluations. We demonstrate that RL such as GRPO can work effectively for small LLMs.
@@ 3.4 Training Configuration
We use Colossal-AI [98] as our training framework with acceleration techniques including optimized parallelism, memory optimization, and heterogeneous training methods. 1) Optimized Parallelism: We combine model, data, and pipeline parallelism to scale training across GPUs. This hybrid parallelism strategy minimizes communication overhead and ensures efficient resource utilization. 2) Memory Optimization: Techniques such as ZeRO redundancy elimination, gradient checkpointing, and mixed-precision training (FP16) are employed to maximize the effective batch size and reduce memory consumption. 3) Heterogeneous Training Methods: the dynamic scheduling and asynchronous communication protocols are leveraged to further boost training throughput. These methods are provided through user-friendly APIs, requiring minimal code modifications. The training cost is around \$160,000.
@@ 6 Evaluation
We conducted comprehensive performance comparisons against leading language models of comparable scale, including Mistral-7B [6], LLaMA 2-7B [71], Gemma-7B [56], and Qwen v2-7B [11]. These models were selected based on their demonstrated excellence within the 7B or 8B category and represent diverse development approaches from various research organizations worldwide. To ensure a robust evaluation, we re-run all benchmarks with the same evaluation pipeline for fair comparisons. Specifically, we use lm-evaluation-harness [111], opencompass [112], Qwen-Math [109], and OLMES [62] for evaluation.
@@ 6.2 Pre-Training Evaluation
We name the initial model as Moxin-7B-Original, which presents the foundation model before fine-tuning on the training data of the evaluation datasets. After subsequent partial fine-tuning of Moxin-7B-Original on the training data of the evaluation datasets, we developed Moxin-7B-Enhanced, enabling direct assessment of how targeted fine-tuning affects model performance. We release our Moxin-7B-Enhanced model as Moxin-7B-Base model.
@@ 6.2.1 Zero-Shot Evaluation
We report the result of base models for zero-shot evaluation in Table 6.2.1. The tasks are listed below. After training with the training data of evaluation tasks, our Moxin-7B-Enhanced can achieve superior performance compared with state-of-the-art (SOTA) baselines. This significant increase from the base model demonstrates the effectiveness of our fine-tuning approach. The improved performance is particularly notable on complex reasoning tasks like PIQA, where the score increased from 78.07% to 82.24%, matching or exceeding several leading models. Consequently, our models emerge as an excellent candidate for real-world applications.
@@
[t] Performance comparison for various models in zero-shot evaluation. Models HellaSwag WinoGrade PIQA ARC-E ARC-C Ave Mistral - 7B 80.39 73.4 82.15 78.28 52.22 73.29 LLaMA 2 - 7B 75.99 69.06 79.11 74.54 46.42 69.02 LLaMA 2 - 13B 79.37 72.22 80.52 77.4 49.06 71.71 LLaMA 3.1 - 8B 78.92 74.19 81.12 81.06 53.67 73.79 Gemma - 7b 80.45 73.72 80.9 79.97 54.1 73.83 Qwen v2 - 7B 78.9 72.38 79.98 74.71 50.09 71.21 Internlm2.5 - 7b 79.14 77.9 80.52 76.16 51.37 73.02 Baichuan2 - 7B 72.25 67.17 77.26 72.98 42.15 66.36 Yi-1.5-9B 77.86 73.01 80.74 79.04 55.03 73.14 DeepSeek - 7B 76.13 69.77 79.76 71.04 44.8 68.3 Moxin - 7B - Original 72.06 66.31 78.07 71.47 48.15 67.21 Moxin - 7B - Enhanced 80.03 75.17 82.24 81.12 58.64 75.44
@@ 6.2.2 Few-Shot Evaluation
Table 6.2.2 presents our zero-shot evaluation results across multiple benchmark tasks. The tasks and their few-show settings are listed below. Thanks to its rigorous and high-quality training corpus, our model demonstrates a remarkable competitive edge in tasks that involve language understanding and knowledge application. Our Moxin-7B-Original achieves superior performance than LLaMA2-7B in this scenario. After training with the training data of evaluation tasks, our Moxin-7B-Enhanced can achieve competitive performance compared with SOTA baselines.
@@
[t] Performance comparison for various models in few-shot evaluation. Model ARC-C Hellaswag MMLU WinoGrade Ave Mistral - 7B 57.59 83.25 62.42 78.77 70.51 LLaMA 3.1 - 8B 54.61 81.95 65.16 77.35 69.77 LLaMA 3 - 8B 55.46 82.09 65.29 77.82 70.17 LLaMA 2 - 7B 49.74 78.94 45.89 74.27 62.21 Qwen 2 - 7B 57.68 80.76 70.42 77.43 71.57 Gemma - 7B 56.48 82.31 63.02 78.3 70.03 Internlm2.5 - 7B 54.78 79.7 68.17 80.9 70.89 Baichuan2 - 7B 47.87 73.89 54.13 70.8 61.67 Yi-1.5-9B 58.36 80.36 69.54 77.53 71.48 Moxin - 7B - Original 53.75 75.46 59.43 70.32 64.74 Moxin - 7B - Enhanced 59.47 83.08 60.97 78.69 70.55
@@ 6.3 Post-Training Evaluation
In our post-training, with Tülu 3, we further fine-tune our base model with SFT and then DPO, leading to our Moxin-7B-SFT model and Moxin-7B-DPO model, respectively. We also adopt the Infinity Instruct dataset [63] to train another version of our instruct model based on our base model, resulting in Moxin-7B-DPO-II. We demonstrate the performance in Table 6.3 for zero-shot evaluation and Table 6.3 for few show evaluation. We also use the OLMES framework from Tulu 3 [62] to evaluate our model across a range of tasks and show the results in Table 6.3. We can observe that our Moxin-7B-DPO model can achieve comparable performance with other SOTA instruct models. We release the Moxin-7B-DPO model as Moxin-7B-Instruct model.
@@ 6.4 CoT Evaluation
Based on our instruct model, we further finetune the model with reasoning data based on SFT and then train using reinforcement learning with GRPO following DeepSeek R1. Starting from the same model after SFT with the OpenThoughts and Open-R1-Math-220k datasets, we use two different RL frameworks, i.e., DeepScaleR [67] and AReal [68], to develop two versions of our reasoning model, respectively, resulting in Moxin-7B-RL-DeepScaleR and Moxin-7B-RL-AReal. We evaluate our reasoning models on competition-level mathematics benchmarks, including AMC 2023, MATH-500, Minerva Math, and OlympiadBench. We report the Pass@1 accuracy. We use Qwen Math’s codebase for evaluation following [109, 110] and show the comparison in Table 6.4. We can observe that the Moxin-7B-RL-DeepScaleR model can achieve outstanding performance compared with baselines, demonstrating the effectiveness of RL for small LLMs such as 7B models. Our Reasoning model trained with DeepScaleR performs better than that trained with AReal. Our reasoning model Moxin-7B-RL-DeepScaleR trained with DeepScaleR [67] is released as Moxin-7B-Reasoning model.
[t] Performance comparison for various models on reasoning evaluation. Models/Datasets MATH 500 AMC Minerva Math OlympiadBench Ave Qwen2.5-Math-7B-Base 52.4 52.5 12.9 16.4 33.55 Qwen2.5-Math-7B-Base + 8K MATH SFT 54.6 22.5 32.7 19.6 32.35 Llama-3.1-70B-Instruct 64.6 30.1 35.3 31.9 40.48 Moxin-7B-RL-AReal 68.6 50 16.9 31.7 41.8 Moxin-7B-RL-DeepScaleR 68 57.5 16.9 30.4 43.2
We demonstrate a list of our models in the paper in Table 6.4. Multiple models are developed in our paper, and we show their names in our releases.
[h] Our developed models and their names in our releases. Developed Models Names in Releases Moxin-7B-Enhanced Moxin-7B-Base Moxin-7B-SFT Moxin-7B-DPO Moxin-7B-Instruct Moxin-7B-DPO-II Moxin-7B-RL-DeepScaleR Moxin-7B-Reasoning Moxin-7B-RL-AReal